OpenStudy (anonymous):

an electric immersion heater has a power rating of 1500 W. If the heater is placed in a liter of water @ a room temperature, how many minutes will it take to bring water to a boil? (assume that there is no heat loss)

OpenStudy (anonymous):

let t minutes be required so Heat energy given by Heater = 1500*t*60 J and so on..

OpenStudy (anonymous):

Heat energy required =Mass*specific heat *change in temperature..

OpenStudy (anonymous):

why 60, is it for seconds ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

you need to first find the mass of 1 ltr of water in kg using density of water..

OpenStudy (anonymous):

Heat energy given by Heater = 1500*t*60 J Heat energy required =1000g (1)(20)

OpenStudy (anonymous):

mass should be in Kg and not in gms and density should be in Kg/m^3 and

OpenStudy (anonymous):

1kg = 1L

OpenStudy (anonymous):

we're using grams instead of kg.

OpenStudy (anonymous):

not necessarily depends on density... you can do it in gms and cm but you need to include J (mechanical equivalent of heat energy i.e conversion between calorie and Joules)

OpenStudy (anonymous):

where will i put the answer in heat energy required?

OpenStudy (anonymous):

so, i'll use 1kg?

OpenStudy (anonymous):

Heat is a form of energy, and therefore is measured in joules. There are other units of heat, though; the most common one is the kilocalorie: One kilocalorie (kcal) is defined as the amount of heat needed to raise the temperature of 1 kg of water by 1 C° (from 14.5°C to 15.5°C).

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

room temp = 20 deg.cent. boiling point = 100 deg.cent. delta in temp rise = 80 deg.cent

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

energy needed to boil the water = 1kg*4186*80 degrees = 4186*80 = 334880 energy given by the heater to boil the water = 1500*t seconds t = 334880 / 1500 = 223.2533 seconds = 3.7 minutes

OpenStudy (anonymous):

1 kCal = 4186 joules

OpenStudy (anonymous):

does the above answer look right?

OpenStudy (anonymous):

why 80 degrees?

OpenStudy (anonymous):

isn't that the rise in tempreture needed to heat the water from room temp 20 degress to 100 degrees boling point

OpenStudy (anonymous):

ahhh,,yeah,,gets

OpenStudy (anonymous):

[mc DeltaT\] 1000grams (1)(80),